# From Plausibility to Verifiability: Risk-Controlled Generative OCR for Vision-Language Models

### 20 Mar 2026 v1

@@ 3.4. Decision Functional: Consensus, Stability, and Operating-Point Control
where $\mathrm{ED}(\cdot,\cdot)$ is Levenshtein edit distance, and define the candidate-centered dispersion
$\Delta(I)=\frac{1}{n(I)}\sum_{k\in\Omega(I)}d\!\left(\tilde{y}_{k},s^{*}(I)\right).$
The resulting evidence summary is
@@
The controller uses a consensus threshold $\tau(m)$ and a fixed dispersion threshold $\kappa$. In the main operating-point family, the strictness index $m$ affects only the minimum required vote fraction, with $\tau(m)$ nondecreasing in $m$, while $\kappa$ is held fixed. Larger $m$ therefore induces a stricter acceptance region through a stronger consensus requirement under the common stability gate $\Delta(I)\leq\kappa$.
The instantiated controller is
@@
Algorithm 1 Multi-view geometric probing and accept–abstain control
1:Crop II, frozen generator $G_{\delta}$, prompt $p$, protocol/controller instantiation $(\pi,m)$
@@
5:  $\tilde{y}_{k}\leftarrow\mathrm{Norm}(G_{\delta}(I_{k},p;\theta))$,  $v_{k}\leftarrow\mathbb{I}[\tilde{y}_{k}\in\mathcal{V}]$
@@
9:Determine the unique mode $s^{*}$ of $\{\tilde{y}_{k}:k\in\Omega\}$
@@ 3.5. Coverage, Risk, and Operating Points
All controlled-system risk metrics are computed on the covered subset. We report mean CER and upper-tail CER (P99) on covered outputs, and define the conditional catastrophic exposure rate as
@@ 4.1. Experimental Setup
Backbones and datasets. We evaluate the controller on three frozen vision-language backbones used as black-box generators: LLaVA-Phi3 (3.8B) (Zhu et al., 2024), Gemma3 (4B) (Team et al., 2025), and GLM-OCR. Experiments use the official test sets of IIIT5K (Mishra et al., 2012) and ICDAR 2013 (Karatzas et al., 2013).
@@ 4.2. Main Results
Figure 3. Qualitative OCR cases under the fixed deployment protocol. Each panel shows the crop, ground truth, the always-accept baseline, and the GRC decision at the default operating point ($m=3$). Cases include over-generation, unsupported substitution, correct acceptance, cross-view instability, a stable-but-wrong residual failure, and a borderline accepted case. The evidence tag reports valid-view count, consensus fraction $q$, and geometric length estimate $L_{\mathrm{geom}}$; for the instability case, it reports dispersion $\Delta$ instead of $L_{\mathrm{geom}}$.
@@
Coverage is reported for the controlled system only. For risk metrics, each entry is shown as Baseline/System, where Baseline is computed on the full test set and System on the covered subset. CER is reported in %; MD@2 in ‰.
Table 1 reports the main deployment-level results under the common fixed protocol with $K=5$ at the moderate operating point $m=3$. These results address the deployment setting introduced in Sec. 1 by quantifying the error risk exposed on the outputs that the system actually returns under a fixed selective contract.
Across all $3\times 2$ model–dataset settings, the Geometric Risk Controller maintains high coverage while substantially reducing exposed risk relative to the always-accept baseline. Relative to baseline full-set exposure, the controlled system consistently lowers mean CER, suppresses upper-tail error, and reduces catastrophic exposure as measured by $\mathrm{Meltdown}@2$. These results align with the intended deployment role of the controller, turning open-ended generative OCR into a selective system that trades a moderate amount of coverage for a substantially safer set of exposed outputs. Because the baseline never abstains, its risk is reported on the full test set; for the controlled system, risk is reported on the covered subset together with coverage. Figure 3 shows representative qualitative outcomes under the fixed protocol, including over-generation, unsupported substitution, instability-driven abstention, and stable-but-wrong residual failures.
@@ 4.3. Approximately Coverage-Matched External Baseline Comparison
For fairness, this baseline uses the same frozen backbone, prompt template, deterministic decoding, and output canonicalization as the always-accept OCR baseline. It operates on a single view and accepts a transcription only when the model’s internal confidence exceeds a calibrated threshold. Confidence is instantiated by the mean token log-probability of the generated transcription, and the threshold is calibrated on a held-out split to target the coverage of our default operating point ($m=3$) for each backbone–dataset setting. Realized test-set coverage is therefore approximately, rather than exactly, matched.
@@
Table 2. Approximately coverage-matched comparison with an external confidence-threshold selective baseline adapted from Srinivasan et al. (2024). For each backbone–dataset setting, the threshold is calibrated on a held-out split to target the coverage of our default operating point ($m=3$). Coverage and mean CER are reported in %; Meltdown@2 (MD@2) is reported in ‰.
@@
4.4. Operating-Point Control
Figure 4. Risk–coverage trajectories under the common fixed protocol for all three backbones on IIIT5K and ICDAR13. Each point is a predefined operating point indexed by $m$; larger $m$ yields lower coverage and lower covered-output risk. Filled markers denote the default operating point $m=3$.
Risk–coverage plots for three backbones: GLM-OCR, Gemma3, and LLaVA-Phi3. The x-axis is coverage and the y-axis is system mean CER on covered outputs. Each subplot shows two curves, one for IIIT5K and one for ICDAR13, with operating points indexed by m. As m increases, the operating point moves toward lower coverage and lower risk. Filled markers indicate the default operating point m=3.
Figure 5. Component effects and comparison to an external confidence-threshold baseline on LLaVA-Phi3 under the fixed protocol (K=5K{=}5, m=3m{=}3). Bars show Meltdown@2; labels show coverage. Full GRC achieves the lowest catastrophic exposure on both datasets while maintaining competitive coverage.
@@ 4.4. Operating-Point Control
A key property of the controller is that the strictness index $m$ does not modify the backbone or protocol; it selects one predefined operating point under a fixed selective contract. In our main instantiation, increasing $m$ raises the required consensus threshold $\tau(m)$ while the dispersion gate remains fixed, making acceptance more conservative and trading coverage for lower exposed risk.
Figure 4 visualizes this behavior across all three backbones on both IIIT5K and ICDAR13. In each case, increasing $m$ moves the system along a consistent risk–coverage frontier toward lower coverage and lower covered-output mean CER. This behavior is consistent with interpreting $m$ as a practical deployment knob, since the operator does not modify the backbone or protocol, but instead selects an operating point from a fixed and auditable family.
@@ 4.5. Ablation Study
Table 3. Component ablation under the common fixed protocol ($K=5$, $m=3$). w/o structural removes the protocol-fixed admissibility screen, w/o consensus removes the cross-view consensus control, and Full GRC applies both. Coverage and CER are reported in %. Meltdown@2 is omitted here and shown in Fig. 5.
Method LLaVA-Phi3 Gemma3 GLM-OCR
[…]
IIIT5K
Baseline 100.0 110.5 100.0 7.4 100.0 6.3
w/o structural 95.3 21.8 98.8 4.6 99.6 5.5
w/o consensus 91.8 11.2 95.8 4.3 96.0 4.5
Full GRC 89.5 8.4 94.9 3.7 95.7 4.0
ICDAR13
Baseline 100.0 147.4 100.0 20.1 100.0 3.3
w/o structural 93.1 31.7 96.1 10.4 99.6 3.3
w/o consensus 93.7 17.4 96.0 8.7 96.5 2.5
Full GRC 88.9 9.3 93.0 3.7 96.2 2.3
We next ablate the two decision channels of the controller under the same fixed protocol used above. Table 3 compares the full system with two partial variants. The w/o structural variant removes the protocol-fixed admissibility screen while retaining cross-view consensus control, whereas w/o consensus removes vote-fraction and dispersion control while retaining structural screening.
The two components play distinct and complementary roles. The w/o consensus variant is particularly effective at filtering obviously degenerate continuations and often delivers strong reductions in catastrophic exposure. The w/o structural variant captures cross-view instability and usually improves over the baseline, but its gains are less uniform, especially on already stronger backbones. Neither partial variant matches the full controller.
Full GRC consistently delivers the strongest overall deployment behavior across models and datasets, achieving the lowest exposed risk in every setting while maintaining high coverage. Taken together, these results show that the controller is not a loose collection of heuristics, but a complementary combination of admissibility screening and cross-view consensus control.
4.6. Operating-Point Endpoints and Practical Trade-offs
Table 4. Representative endpoint operating points under the fixed protocol ($K=5$). Coverage (Cov.) is reported in %, and Meltdown@2 (MD@2) is reported in ‰.
@@ 4.6. Operating-Point Endpoints and Practical Trade-offs
Figure 4 shows that predefined operating points trace a risk–coverage frontier under a fixed protocol. Because deployment is typically selected from a small set of discrete operating-point options, Table 4 reports three representative settings, namely the always-accept baseline, a permissive setting ($m=1$), and a conservative setting ($m=5$). Stricter operating points reduce catastrophic exposure at the cost of lower coverage. This separation is most pronounced on LLaVA-Phi3 and remains consistent on Gemma3 and GLM-OCR. The default setting $m=3$ used in Sec. 4.2 lies between these endpoints and provides a practical balance between availability and exposed-output reliability.
Table 5 reports a compact query-budget check on LLaVA-Phi3. The single-pass always-accept baseline is included for reference, while $K\in\{3,5,7\}$ varies the evidence budget within the multi-view controller. Reducing to $K=3$ slightly lowers catastrophic exposure but also lowers coverage, while increasing to $K=7$ adds cost with little gain over $K=5$. This supports $K=5$ as a practical default.
Table 5. Compact query-budget check for LLaVA-Phi3 at $m=3$. Relative time is normalized by the single-pass baseline; coverage is reported in %, and Meltdown@2 in ‰.

### 15 Apr 2026 v3

@@ 3.4. Decision Functional and Operating-Point Control
where $d(a,b)$ is the bounded normalized edit distance between strings $a$ and $b$, $\mathrm{ED}(a,b)$ is their Levenshtein edit distance, and $|a|$ denotes the length of string $a$. Using $d(\cdot,\cdot)$, we define the candidate-centered dispersion
$\Delta(I)=\frac{1}{n(I)}\sum_{k\in\Omega(I)}d\!\left(\tilde{y}_{k},s^{*}(I)\right),$
which measures how tightly the valid outputs cluster around the candidate. The resulting evidence summary is
@@
The controller uses two thresholds: a consensus threshold $\tau(m)$ and a dispersion threshold $\kappa$. Here, $m$ is the strictness index of the operating point, and $\tau(m)$ is nondecreasing in $m$. In the main operating-point family, changing $m$ affects only the minimum required vote fraction, while $\kappa$ is held fixed. Larger $m$ therefore induces a stricter acceptance region through a stronger consensus requirement under the common stability gate $\Delta(I)\leq\kappa$.
The instantiated controller under protocol $\pi$ and operating point $m$ is
@@
where $\bot$ denotes abstention.
Figure 2 provides an intuitive view of this accept/abstain rule under the fixed K=5K{=}5 protocol at the default operating point m=3m{=}3. In the top row, fragmented support leads to abstention because the vote fraction $q(I)$ is too low. In the middle row, a majority candidate still leads to abstention because the valid outputs remain too dispersed, yielding a large $\Delta(I)$. In the bottom row, a borderline majority is accepted because the valid views remain tightly clustered, so both the consensus and stability criteria are satisfied.
Figure 2. Representative accept/abstain evidence patterns under the fixed K=5K{=}5 protocol at the default operating point m=3m{=}3. Fragmented valid outputs cause abstention, a majority can still be rejected under high dispersion, and borderline consensus is accepted when the valid views remain tightly clustered.
Algorithm 1 summarizes the full inference-time execution of GRC under a fixed protocol/controller pair $(\pi,m)$. Starting from input crop II, it generates $K$ protocol-specified views, queries the frozen generator on each view, canonicalizes the returned strings, applies structural screening, and then evaluates the operating-point rule using $q$ and $\Delta$.
Algorithm 1 Inference-time execution of the Geometric Risk Controller under a fixed protocol
1:Crop II; frozen generator $G_{\theta,\delta}$; prompt $p$; protocol/controller pair $(\pi,m)$
@@
5:  $\tilde{y}_{k}\leftarrow\mathrm{Norm}(G_{\theta,\delta}(I_{k},p))$,  $v_{k}\leftarrow\mathbb{I}[\tilde{y}_{k}\in\mathcal{V}]$
@@
9:Let $s^{*}$ be the unique mode of $\{\tilde{y}_{k}:k\in\Omega\}$
@@
Algorithm 1 is an executable summary of Eq. (1). The branch $|\Omega|<K_{\min}$ implements the minimum-evidence gate induced by structural screening, while the final decision branch applies the operating-point rule through $\tau(m)$ and $\kappa$.
@@ 3.5. Coverage, Risk, and Operating Points
Let $y^{\mathrm{gt}}$ denote the ground-truth transcription. All controlled-system risk metrics are computed on the covered subset. We report mean character error rate (CER), 99th-percentile CER (P99), and the conditional catastrophic exposure rate
@@ 4.1. Experimental Setup
Backbones and datasets. We evaluate the controller on three frozen vision-language backbones used as black-box generators: LLaVA-Phi3 (3.8B) (Zhu et al., 2024), Gemma3 (4B) (et al., 2025), and GLM-OCR (Duan et al., 2026). Experiments use the official test sets of IIIT5K (Mishra et al., 2012) and ICDAR 2013 (Karatzas et al., 2013).
@@ 4.2. Main Results
Figure 3. Qualitative OCR cases under the fixed deployment protocol. Each panel shows the crop, ground truth, the always-accept baseline, and the GRC decision at the default operating point $m=3$. Cases include over-generation, unsupported substitution, correct acceptance, cross-view instability, a stable-but-wrong residual failure, and a borderline accepted case. The evidence tag reports valid-view count, consensus fraction $q$, and geometric length estimate $L_{\mathrm{geom}}$; for the instability case, it reports dispersion $\Delta$ instead of $L_{\mathrm{geom}}$.
@@
For each metric, entries are shown as Baseline/System. For coverage, the baseline is always 100.0 because the always-accept system never abstains. For risk metrics, Baseline is computed on the full test set and System on the covered subset. CER is reported in %, MD@2 in ‰.
Table 1 reports the main deployment-level results under the common fixed protocol with $K=5$ at the moderate operating point $m=3$. These results address the deployment setting introduced in Sec. 4.1 by quantifying the error risk exposed on the outputs that the system actually returns under a fixed selective contract.
Across all $3\times 2$ model–dataset settings, the geometric risk controller maintains high coverage while substantially reducing exposed risk relative to the always-accept baseline. Relative to baseline full-set exposure, the controlled system consistently lowers mean CER, suppresses upper-tail error, and reduces catastrophic exposure as measured by $\mathrm{Meltdown}@2$. These results align with the intended deployment role of the controller, turning open-ended generative OCR into a selective system that trades a moderate amount of coverage for a substantially safer set of exposed outputs. Figure 3 shows representative qualitative outcomes under the fixed protocol, including over-generation, unsupported substitution, instability-driven abstention, and stable-but-wrong residual failures.
@@ 4.3. Approximately Coverage-Matched External Baseline Comparison
For fairness, this baseline uses the same frozen backbone, prompt template, deterministic decoding, and output canonicalization as the always-accept OCR baseline. It operates on a single view and accepts a transcription only when the model’s internal confidence exceeds a calibrated threshold. Confidence is instantiated by the mean token log-probability of the generated transcription, and the threshold is calibrated on a held-out split to target the coverage of our default operating point $m=3$ for each backbone–dataset setting. Realized test-set coverage is therefore approximately, rather than exactly, matched.
@@
Table 2. Approximately coverage-matched comparison with an external confidence-threshold selective baseline adapted from Srinivasan et al. (2024). For each backbone–dataset setting, the threshold is calibrated on a held-out split to target the coverage of our default operating point $m=3$. Coverage and mean CER are reported in %, Meltdown@2 (MD@2) is reported in ‰.
@@
$4.$ Operating-Point Control
Figure 4. Risk–coverage trajectories under the common fixed protocol for all three backbones on IIIT5K and ICDAR13. Each subfigure corresponds to one backbone: (a) GLM-OCR, (b) Gemma3, and (c) LLaVA-Phi3. Each point is a predefined operating point indexed by $m$ while larger $m$ yields lower coverage and lower covered-output risk. Filled markers denote the default operating point $m=3$.
A figure with three horizontally arranged subplots showing risk–coverage trajectories for three OCR backbones: (a) GLM-OCR, (b) Gemma3, and (c) LLaVA-Phi3. In each subplot, the horizontal axis is coverage in percent and the vertical axis is mean character error rate on covered outputs in percent. Two dataset-specific curves are shown, one for ICDAR13 and one for IIIT5K. Each curve consists of predefined operating points indexed by m, with labels marking values such as m=1, m=3, and m=5. Filled markers denote the default operating point m=3. Across all three models, larger m corresponds to lower coverage and lower covered-output risk, with the strongest separation visible for LLaVA-Phi3. A shared legend identifies the two datasets.
Figure 5. Component effects and comparison to an external confidence-threshold baseline on LLaVA-Phi3 under the fixed protocol K=5K{=}5, m=3m{=}3. Bars show Meltdown@2, and labels show coverage. Both structural screening and cross-view consensus reduce catastrophic exposure, while Full GRC achieves the lowest Meltdown@2 on both datasets with competitive coverage.
@@
A key property of the controller is that the strictness index $m$ does not modify the backbone or protocol. Instead, it selects a predefined operating point under a fixed selective contract. In our main instantiation, increasing $m$ raises the required consensus threshold $\tau(m)$ while the dispersion gate remains fixed, making acceptance more conservative and trading coverage for lower exposed risk. This design makes $m$ a practical deployment knob: the operator does not alter model parameters or inference procedures, but instead selects from a fixed and auditable family of operating points.
Figure 4 visualizes the full risk–coverage trajectories induced by these predefined operating points under the fixed protocol, while Table 3 reports exact numbers for three representative deployment settings, namely the always-accept baseline, a permissive setting $m=1$, and a conservative setting $m=5$. Across all three backbones, increasing strictness moves the system toward lower coverage and lower exposed risk. The figure emphasizes the global trade-off frontier, whereas the table highlights deployment-relevant endpoint statistics, especially catastrophic exposure measured by MD@2. This separation is most pronounced on LLaVA-Phi3 and remains consistent on Gemma3 and GLM-OCR. The default setting $m=3$ used in Sec. 4.2 lies between these endpoints and provides a practical balance between availability and exposed-output reliability.
@@ 4.5. Ablation Study
We next ablate the two decision channels of the controller under the same fixed protocol used above. Figure 5 compares the full system with two partial variants and an external confidence-threshold baseline on LLaVA-Phi3. The w/o structural variant removes the protocol-fixed admissibility screen while retaining cross-view consensus control, whereas w/o consensus removes vote-fraction and dispersion control while retaining structural screening.
The two components play distinct and complementary roles. The w/o structural variant reduces catastrophic exposure relative to the always-accept baseline, showing that cross-view consensus alone already provides meaningful risk control. The w/o consensus variant also improves substantially over the baseline, indicating that structural screening is effective at filtering degenerate continuations. However, neither partial variant matches the full controller.
Full GRC achieves the lowest Meltdown@2 on both IIIT5K and ICDAR13 while maintaining competitive coverage. The external confidence-threshold baseline reduces catastrophic exposure compared with the always-accept baseline, but remains clearly worse than Full GRC on both datasets. Taken together, these results show that the controller is not a loose collection of heuristics, but a complementary combination of admissibility screening and cross-view consensus control.
4.6. Effect of Query Budget
Table 3. Representative endpoint operating points under the fixed protocol $K=5$. Coverage (Cov.) is reported in %, and Meltdown@2 (MD@2) is reported in ‰.
@@ 4.6. Effect of Query Budget
Table 4 reports a compact query-budget check on LLaVA-Phi3. The single-pass always-accept baseline is included for reference, while $K\in\{3,5,7\}$ varies the evidence budget within the multi-view controller under the same fixed operating protocol. Reducing the number of views to $K=3$ slightly lowers catastrophic exposure, but also lowers coverage, reflecting a more conservative controller with reduced availability. By contrast, increasing the budget to $K=7$ introduces additional query cost with little improvement over $K=5$. Taken together, these results indicate that $K=5$ already captures most of the benefit of multi-view evidence aggregation, making it a practical default for deployment.
Table 4. Compact query-budget check for LLaVA-Phi3 at $m=3$. Relative time is normalized by the single-pass baseline and coverage is reported in %, and Meltdown@2 in ‰.